A car traveling at 45 mph. The tires are 30 in in diameter. What is the angular velocity of the tire?
i got v=rw
r= 15in(1ft/12in)(1mi/5280ft) so r= 2.37E-4
V= 45mph
w= v/r
45mi/2.37E-4(mi)(hr)
which leaves 189873.418/hr(hr/3600sec)
so w= 52.743/sec
i dont know how to show work to get rad/sec can i just times it by 1 rad? or was the 45/2.37E-4 the rad?

Answer by Alan3354(31538)

A car traveling at 45 mph. The tires are 30 in in diameter. What is the angular velocity of the tire?
Use feet, not miles.
45 mph = 66 ft/sec
------------------
One rev of the tire is 30*pi inches = 2.5pi feet
66 ft/sec / 2.5pi ft/rev = 26.4/pi rev/sec
1 rev = 2pi radians
26.4/pi * 2pi = 52.8 radians/sec
----------------------------------
52.743 is the right answer, just with roundoff errors
==================================
i got v=rw
r= 15in(1ft/12in)(1mi/5280ft) so r= 2.37E-4
V= 45mph
w= v/r
45mi/2.37E-4(mi)(hr)
which leaves 189873.418/hr(hr/3600sec)
so w= 52.743/sec
i dont know how to show work to get rad/sec can i just times it by 1 rad? or was the 45/2.37E-4 the rad?